Suppose that you are filling an aquarium with a hose of radius 2 cm. You measure the volume flow rate and find that it is 5 L per minute. If all other factors remain constant, what would the flow rate be if you switched to a hose with a radius of 5 cm? Show the steps taken to find your answer.
Answer #1

Volumetric flow rate is given as:

Q = Area of cross section x Velocity

Q = A x V

Now Q1 = A1 x V1 ............(i)

Q2 = A2 x V2...................(ii)

Dividing (ii) by (i)

2 Qi A1Vi

Since all other factors are kept constant hence V1 = V2

\frac{Q_2}{Q_1} = \frac{A_2}{A_1}

5-2 π-π

5- Q2 = Q1

52 Q2 = 5 × 2 = 31.25L/min man